I need help factoring this problem: x^2 - 11xy + 24y^2
the ^ sign is meant to represent the power sign.
Thank you

Answer by KnightOwlTutor(293)

First we have to determine what two numbers give a product of 24
12,2
8,3
6,4
24,1

We know that both of the numbers are - because the final term is +
Let us find two of the numbers when added together=-11
-12-2=-14
-8-3=-11
The answer is
(x-8y)(x-3y)